The following table shows betas for several companies. Calculate each stock’s expected rate of return using the CAPM. Assume the risk-free rate of interest is 6%. Use a 8% risk premium for the market portfolio. (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places.)

Company Beta Cost of Capital
Caterpillar 1.81 %
Apple 1.45 %
Johnson & Johnson 0.64 %
Consolidated Edison 0.36 %
Answer #1

Risk-free rate = RF = 6%

Market risk premium = (RM - RF) = 8%

CAPM Equation

Expected return on an investemnt i is given by:

E[Ri] = RF + βi*(RM - RF)

where, RF = Risk-free rate, βi = beta of the investment i, (RM - RF) = market-risk premium

Caterpillar

Beta of caterpillar = βC = 1.81

Expected return on Caterpillar = E[RC] = RF + βA*(RM - RF) = 6% + 1.81*8% = 20.48%

Apple

Beta of Apple = βA = 1.45

Expected return on Apple = E[RA] = RF + βA*(RM - RF) = 6% + 1.45*8% = 17.6%

Johnson & Johnson

Beta of Johnson & Johnson = βJ = 0.64

Expected return on Caterpillar = E[RJ] = RF + βJ*(RM - RF) = 6% + 0.64*8% = 11.12%

Consolidated Edison

Beta of Edison = βE = 0.36

Expected return on Edison = E[RE] = RF + βE*(RM - RF) = 6% + 0.36*8% = 8.88%

Answer (in %)

Company Cost of Capital
Caterpillar 20.48
Apple 17.60
Johnson & Johnson 11.12
Consolidated Edison 8.88